Solve the following logarithmic equation; log(2x-1) = log(x+3) + log3
Answer #1

Sol. log (2x - 1) = log (2+3) + leg 3 [ loga & logb = log ab ] log (22-1) = log [ 3(3+3)] dog (2n-1) - log [3ntg J so [loga-Verifying Solution dog (2n-1) = = dog [2 (-10)-1] log (-21) undefined log (2+3) + log3 a dog [3n+g] = d.g [3(-10) to - dog [-